OpenStudy (anonymous):

If the heat of combustion for a specific compound is -1340.0 kJ/mol and its molar mass is 91.17 g/mol, how many grams of this compound must you burn to release 149.30 kJ of heat?

OpenStudy (abb0t):

\(\sf \color{}{ \frac{-13400.0~kJ}{mol} \times \frac{mol}{91.17~g} \rightarrow 149.39 ~kj \times \frac{g}{some~number~kJ} }\)